# CREAM: Continual Retrieval on Dynamic Streaming Corpora with Adaptive Soft Memory

### v1

@@ Abstract.
Information retrieval (IR) in dynamic data streams is emerging as a challenging task, as shifts in data distribution degrade the performance of AI-powered IR systems. To mitigate this issue, memory-based continual learning has been widely adopted for IR. However, existing methods rely on a fixed set of queries with ground-truth relevant documents, which limits generalization to unseen queries and documents, making them impractical for real-world applications. To enable more effective learning with unseen topics of a new corpus without ground-truth labels, we propose CREAM, a self-supervised framework for memory-based continual retrieval. CREAM captures the evolving semantics of streaming queries and documents into dynamically structured soft memory and leverages it to adapt to both seen and unseen topics in an unsupervised setting. We realize this through three key techniques: fine-grained similarity estimation, regularized cluster prototyping, and stratified coreset sampling. Experiments on two benchmark datasets demonstrate that CREAM exhibits superior adaptability and retrieval accuracy, outperforming the strongest method in a label-free setting by 27.79% in Success@5 and 44.5% in Recall@10 on average, and achieving performance comparable to or even exceeding that of supervised methods.
@@ 1.1. Background
Information retrieval (IR) in online environments, powering real-time retrieval-augmented generation (lewis2020retrieval) and agentic context engineering (zhang2025agentic), is emerging as a key technology for various downstream applications. For example, in a real-time news summarization system (pdsum), a query “What are the current issues in the global supply chain?” would require retrieving relevant articles covering current events, such as geopolitical conflicts or new tariff policies, at the time of the query. Identifying query-relevant documents in streaming corpora with dynamically evolving topics is challenging, as pretrained encoders become outdated under domain shifts, impairing relevance matching performance. This challenge is especially critical in real-world IR systems requiring timely and accurate responsiveness, which is more pronounced in emerging agentic AI frameworks that facilitate real-time decision-making (agentic_ai).
Specifically, consider a scenario where the topics in streaming corpora gradually shift from the medical domain to the business domain. In the query “Has the agent been approved?”, the term agent typically refers to a drug in the medical domain, whereas it denotes a person or agency in the business context. If the IR system has not sufficiently adapted to the new domain, it may return irrelevant medical documents, potentially leading to an incorrect answer such as “The FDA approved the therapeutic agent.” In contrast, if a system could rapidly adapt to the emerging domain while retaining relevant knowledge from the previous domain, it is capable of retrieving appropriate business-related documents and producing a more appropriate answer such as “The licensing board approved the real estate agent’s application.
@@ 1.2. Existing Efforts
In a typical IR system, an encoder is optimized to enhance the semantic similarity between query-document pairs labeled as relevant, with these labels obtained through human annotation. As illustrated in Figure 1, when the distribution of queries and documents evolves over time with diverse topics, matching relevant pairs becomes increasingly difficult unless the encoder is continually updated to reflect the evolving corpora. In practice, web-scale corpora in typical IR systems involve a continuous influx of documents and queries, making retraining on all past data for training computationally inefficient and often infeasible. The naive incremental update of the encoder, however, suffers from catastrophic forgetting, a well-known issue in deep learning where previously acquired knowledge is overwritten by new information (catastrophic_forgetting). To address this challenge, existing continual retrieval methods adopt memory-based continual learning strategies (er; mir; gss; ocs; l2r; clever) to acquire new knowledge without forgetting the old one.
As shown in the upper part of Figure 1, existing methods with memory-based continual learning strategies employ dedicated storage for a fixed set of given queries and their corresponding ground-truth documents, which can be referred to as hard memory. Query-document pairs are sampled from hard memory to update the encoder, while new documents relevant to the predefined queries are added from the streaming corpora. This approach has proven effective in scenarios where the topical distribution of the corpus remains relatively stable and consistent with the predefined query set (l2r). However, simply reusing shift-unaware, predefined query-document pairs stored in hard memory for continual training can cause the encoder to learn information that is less relevant to the current topic distribution, leading to poor adaptation to distributional changes. Moreover, in real-time applications (ustory), human-curated supervision from a set of predefined queries with ground-truth documents is not always available in a timely manner (human_ai), making the hard memory strategy impractical. As a result, such methods fail to support effective retrieval on newly emerging topics, and may degrade performance on the initial or ongoing topics.
@@ 1.3. Main Idea and Contributions
  • Fine-grained similarity estimation: In the absence of external supervision from labeled query-document pairs, a simple similarity estimation based on conventional single-vector representations is insufficient to capture the complex and evolving semantics of corpora. This limitation is practically significant, as self-supervision with noisy relevance signals can lead to critical degradation of the encoder. Thus, we fully exploit the entire token-level information to compute fine-grained semantic similarities both for memory construction and retrieval, inspired by the contextualized late interaction (colbertv2). This enables the encoder to robustly adapt to subtle contextual shifts and emerging topics, even without direct supervision signals.
@@ 2.1. Information Retrieval
Traditional information retrieval approaches are often categorized into sparse, dense, and generative retrieval. Sparse retrieval (SR), such as BM25 (bm25), relies on term frequency and inverse document frequency to compute relevance scores based on exact token matches. While efficient and interpretable, these methods suffer from key limitations: they rely heavily on exact string matches and fail to capture the contextual nuance of semantically similar expressions. As a result, they often underperform in settings where lexical variation or richer semantic understanding is required.
Dense retrieval (DR) addresses these issues by encoding queries and documents into dense vector representations using an encoder. Cross-encoders (bert_reranker) jointly encode the concatenated query and document, employing a final linear layer to map the aggregate sequence representation to a scalar similarity score. Although highly effective, this approach requires pairwise computation between all query-document pairs, which is computationally expensive. In contrast, dual-encoders (dpr) independently encode queries and documents, enabling fast retrieval via cosine similarity on precomputed embeddings. While this method significantly reduces inference time, it may struggle with fine-grained matching due to the lack of deep interaction between the query and the document. To address this trade-off, ColBERT (khattab2020colbert) has introduced late interaction, which balances efficiency and efficacy by delaying fine-grained interactions until the retrieval stage.
With the rise of generative models, generative retrieval (GR) has emerged as a new concept in IR. Differentiable Search Index (DSI) (dsi) first proposed the concept of GR which the model generates document identifiers auto-regressively given a query. They showed that larger models achieved greater performance gains, but also that the gains diminished on larger corpora. While DR and GR leverage language models and share the high-level goal of retrieving relevant documents given a query, they formulate the retrieval problem differently. DR focuses on accurately computing and ranking the similarity between queries and documents, whereas GR aims to generate the correct document identifier for a given query, making direct comparison between them inadequate.
@@ 2.2. Continual Learning on IR
In a continual retrieval setting, where new documents and queries continuously arrive, a retrieval model needs to be repeatedly updated to return the most relevant documents for a given query in the latest context. Naively updating the model with the new data can lead to catastrophic forgetting, particularly for DR and GR models, where the models overwrite previously acquired knowledge and lose their generalizability on earlier data. Among the various continual learning strategies that can be applied to mitigate this issue, Memory replay (er; mir; gss; ocs) has been widely adopted due to its simplicity and effectiveness, especially when combined with a contrastive learning objective (scstory). These memory-based approaches typically maintain an external memory initialized with a set of queries and their corresponding relevant document pairs (i.e., positive samples). As new documents arrive, the encoder is updated to make each query closer to the positive samples while pushing it away from irrelevant documents (i.e., negative samples). Recent works such as L2R (l2r) and CLEVER (clever) represent state-of-the-art frameworks for DR- and GR-based continual retrieval, respectively. L2R defines negativity and diversity metrics to mitigate false negatives during negative sampling. CLEVER leverages external memory for a pseudo-query generator with a reconstruction-based objective, a regularization-based objective, and dynamic indexing via Incremental Product Quantization (IPQ). However, both approaches present practical limitations in their supervision assumption: L2R requires a fixed set of queries along with ground-truth documents for training, and CLEVER also depends on a large number of positive query-document pairs for effective model initialization.
@@ 4.1. Retrieval
An effective retrieval system requires capturing the subtle contextual semantics between queries and documents, which is more pronounced in a label-free setting for continual retrieval. To minimize the loss of information from token embeddings of queries and documents, we aim to preserve token-level granularity in relevance estimation. Specifically, instead of relying on a single embedding derived from mean pooling or the [CLS] token, we adopt a token-level similarity approach inspired by ColBERT’s late interaction (colbertv2) to preserve the semantic granularity of individual tokens. Unlike ColBERT, however, we do not modify the encoder architecture or use special tokens to maintain simplicity and efficiency.
@@ 4.2. Memory Update
A benefit function is defined to trade off the gain in accuracy against the computational cost. The Johnson-Lindenstrauss (JL) lemma (JL_lemma) provides a cost model in terms of distortion $\varepsilon$, while the approximation quality imposes a lower bound on $\varepsilon$. Maximizing the benefit within this feasible range yields the optimal distortion rate $\varepsilon^{*}=\frac{1}{3\sqrt{e}}$. Using the JL lemma, we derive the minimal bit size for LSH at $\varepsilon^{*}$. See Appendix A.1 for the full proof. ∎
For example, in our evaluation setting with the LOTTE dataset (colbertv2), each session includes approximately 2,430 queries (with an average of 9 tokens) and 500,000 documents (with an average of 159 tokens), resulting in up to 80 million token embeddings per session, each with 768 dimensions by BERT (bert). Then, according to Theorem 4.1, to maintain an acceptable distortion rate of $\varepsilon=\frac{1}{3\sqrt{e}}\approx 0.2$, the sufficient number of RP-LSH bits is given by:
@@
CREAM summarizes each cluster by the compact triplet $\langle N,LS,SS\rangle$, similar to the cluster feature vector in BIRCH (birch_cfv). In CREAM, however, we track only the distance summaries, which are even more compact than the original embedding summaries in BIRCH. Specifically, $N$ denotes the number of instances in the cluster, $LS$ is the linear sum of distances to the prototype, and $SS$ is the sum of squared distances to the prototype. This cluster summary is sufficient to compute key cluster statistics, such as the mean and standard deviation of distances to the prototype, and also allows efficient incremental updates in an additive manner.
@@ 4.3. Encoder Update
To construct a training sample that effectively reflects the entire knowledge space in the memory, we propose a query sampling strategy by stratified sampling and coreset selection (coreset). A relevant approach, topic-aware sampling (topic_aware_sampling), has shown effectiveness in a static IR setting. While it randomly samples queries from a limited subset of clusters, CREAM considers the full cluster distribution and explicitly selects an optimal query set by maximizing coverage over the entire cluster space.
@@ 5.1. Experiment Setup
Datasets. To effectively model the dynamics of evolving data distributions, we conduct experiments on two real-world benchmark datasets widely used for continual retrieval (l2r; continual_msmarco): LoTTE (colbertv2) and MSMARCO (modified_msmarco) with the queries clustered by topic from the original dataset (msmarco). LoTTE includes five domains (writing, recreation, science, technology, and lifestyle) from StackExchange questions and answers. MSMARCO includes five domains (Names and Public Figures, Dated Events, Pricing/Units, Medical Treatments and Biology/Physics.) from Bing questions and answers. Each dataset is simulated in 10 streaming sessions with partially overlapping topics following the convention in continual learning (clsurvey; dark_bbcl). Based on the two evaluation protocols formalized in Section 3, the training and evaluation query sets are disjoint, and the training and evaluation document sets are either separated or shared. Details of datasets and sessions are given in Appendix A.4.
@@ A.5. Implementation Details
We use the BM25 [bm25] implementation from the Okapi library with $k_{1}=1.5$, $b=0.75$, and $\epsilon=0.25$. For ColBERT [colbertv2], we use the implementation provided in the official L2R codebase. Following the original paper, we set the output dimension of the linear projection layer in the model to 128. ColBERT+ performs incremental learning using negatives sampled from BM25-retrieved (but unannotated) documents. We use the official L2R implementations of Experience Replay (ER) [er], Maximally Interfered Retrieval (MIR) [mir], and Gradient-based Sample Selection (GSS) [gss]. Online Coreset Selection (OCS) [ocs] is implemented based on the L2R codebase, with $\alpha=1.0$, $\beta=1.0$, and $\gamma=1000.0$, following the original paper and code. For L2R [l2r], we use the official implementation with $\alpha=0.6$ and $\beta=0.4$, as configured in the code.
@@ A.8. LSH Bit Size for MSMARCO
Each session includes approximately 2,430 queries (with an average of 32 tokens) and 30,000 documents (with an average of 256 tokens), resulting in up to 8 million token embeddings per session, each with 768 dimensions by BERT [bert]. Then, according to Theorem 4.1, to maintain an acceptable distortion rate of $\varepsilon=\frac{1}{3\sqrt{e}}\approx 0.2$, the minimum number of RP-LSH bits required is:

### v2

@@ Abstract.
Information retrieval (IR) in dynamic data streams is a crucial task, as shifts in data distribution degrade the performance of AI-powered IR systems. To mitigate this issue, memory-based continual learning has been widely adopted for IR. However, existing methods rely on a fixed set of queries with ground-truth documents, which limits generalization to unseen data, making them impractical for real-world applications. To enable more effective learning with unseen topics of a new corpus without ground-truth labels, we propose CREAM, a self-supervised framework for memory-based continual retrieval. CREAM captures the evolving semantics of streaming queries and documents into dynamically structured soft memory and leverages it to adapt to both seen and unseen topics in an unsupervised setting. We realize this through three key techniques: fine-grained similarity estimation, regularized cluster prototyping, and stratified coreset sampling. Experiments on two benchmark datasets demonstrate that CREAM exhibits superior adaptability and retrieval accuracy, outperforming the strongest method in a label-free setting by 27.79% in Success@5 and 44.5% in Recall@10 on average, and achieving performance comparable to or even exceeding that of supervised methods.
@@ 1.1. Background
Information retrieval (IR) in online environments, powering real-time retrieval-augmented generation (Lewis et al., 2020) and agentic context engineering (Zhang et al., 2025), is emerging as a key technology for various downstream applications. For example, in a real-time news summarization system (Yoon et al., 2023a), a query “What are the current issues in the global supply chain?” would require retrieving relevant articles covering current events, such as geopolitical conflicts or new tariff policies, at the time of the query. Identifying relevant documents with dynamically evolving topics is challenging, as pretrained retrieval models become outdated under domain shifts. This challenge is especially critical in real-world IR systems requiring timely and accurate responses, which is more pronounced in emerging agentic AI frameworks that facilitate real-time decision-making (Sapkota et al., 2025).
Specifically, consider a scenario where topics gradually shift from the medical domain to the business domain. In a query “Has the agent been approved?”, the term agent typically refers to a drug in the medical domain, whereas it denotes a person or agency in the business context. If an IR system has not adapted to the new domain, it may return irrelevant medical documents, potentially leading to an incorrect answer such as “The FDA approved the therapeutic agent.” In contrast, if a system rapidly adapts to the emerging domain while retaining relevant knowledge from the previous domain, it is capable of retrieving appropriate business-related documents and producing a more appropriate answer, such as “The licensing board approved the real estate agent’s application.
@@ 1.2. Existing Efforts
In a typical IR system, an encoder is optimized to enhance the semantic similarity between query-document pairs labeled as relevant, with these labels obtained through human annotation. As illustrated in Figure 1, when the distribution of queries and documents evolves over time with diverse topics, matching relevant pairs becomes increasingly difficult unless the encoder is continually updated to reflect the evolving corpora. In practice, web-scale corpora in typical IR systems involve a continuous influx of documents and queries, making retraining on all past data for training computationally inefficient and often infeasible. The naive incremental update of the encoder, however, suffers from catastrophic forgetting, a well-known issue in deep learning where previously acquired knowledge is overwritten by new information (French, 1999). To address this challenge, existing continual retrieval methods adopt memory-based continual learning strategies (Rolnick et al., 2019; Aljundi et al., 2019a, b; Yoon et al., 2021; Cai et al., 2023; Chen et al., 2023) to acquire new knowledge without forgetting the old one.
As shown in the upper part of Figure 1, existing methods with memory-based continual learning strategies employ dedicated storage for a fixed set of given queries and their corresponding ground-truth documents, which can be referred to as hard memory. Query-document pairs are sampled from hard memory to update the encoder, while new documents relevant to the predefined queries are added from the streaming corpora. This approach has proven effective in scenarios where the topical distribution of the corpus remains relatively stable and consistent with the predefined query set (Cai et al., 2023). However, simply reusing shift-unaware, predefined query-document pairs stored in hard memory for continual training can cause the encoder to learn information that is less relevant to the current topic distribution, leading to poor adaptation to distributional changes. Moreover, in real-time applications (Yoon et al., 2023b), human-curated supervision from a set of fixed queries with ground-truth documents is not always available in a timely manner (Monarch, 2021), making the hard memory strategy impractical. As a result, such methods fail to support effective retrieval on newly emerging topics, and may degrade performance on the initial or ongoing topics.
@@ 1.3. Main Idea and Contributions
  • Fine-grained similarity estimation: In the absence of external supervision from labeled query-document pairs, a simple similarity estimation based on conventional single-vector representations is insufficient to capture the complex and evolving semantics of corpora. This limitation is practically significant, as self-supervision with noisy relevance signals can lead to critical degradation of the encoder. Thus, we fully exploit the entire token-level information to compute fine-grained semantic similarities both for memory construction and retrieval, inspired by the contextualized late interaction (Santhanam et al., 2022). This enables the encoder to robustly adapt to subtle contextual shifts and emerging topics, even without direct supervision signals.
@@ 2.1. Information Retrieval
Traditional information retrieval approaches are often categorized into sparse, dense, and generative retrieval. Sparse retrieval (SR), such as BM25 (Robertson et al., 2009), relies on term frequency and inverse document frequency to compute relevance scores based on exact token matches. While efficient and interpretable, these methods suffer from key limitations: they rely heavily on exact string matches and fail to capture the contextual nuance of semantically similar expressions. As a result, they often underperform in settings where lexical variation or richer semantic understanding is required.
Dense retrieval (DR) addresses these issues by encoding queries and documents into dense vector representations using an encoder. Cross-encoders (Gao et al., 2021) jointly encode the concatenated query and document, employing a final linear layer to map the aggregate sequence representation to a scalar similarity score. Although highly effective, this approach requires pairwise computation between all query-document pairs, which is computationally expensive. In contrast, dual-encoders (Karpukhin et al., 2020) independently encode queries and documents, enabling fast retrieval via cosine similarity on precomputed embeddings. While this method significantly reduces inference time, it may struggle with fine-grained matching due to the lack of deep interaction between the query and the document. To address this trade-off, ColBERT (Khattab and Zaharia, 2020) has introduced late interaction, which balances efficiency and efficacy by delaying fine-grained interactions until the retrieval stage.
With the rise of generative models, generative retrieval (GR) has emerged as a new concept in IR. Differentiable Search Index (DSI) (Tay et al., 2022) first proposed the concept of GR which the model generates document identifiers auto-regressively given a query. They showed that larger models achieved greater performance gains, but also that the gains diminished on larger corpora. While DR and GR leverage language models and share the high-level goal of retrieving relevant documents given a query, they formulate the retrieval problem differently. DR focuses on accurately computing and ranking the similarity between queries and documents, whereas GR aims to generate the correct document identifier for a given query, making direct comparison between them inadequate.
@@ 2.2. Continual Learning on IR
In a continual retrieval setting, where new documents and queries continuously arrive, a retrieval model needs to be repeatedly updated to return the most relevant documents for a given query in the latest context. Naively updating the model with the new data can lead to catastrophic forgetting, particularly for DR and GR models, where the models overwrite previously acquired knowledge and lose their generalizability on earlier data. Among the various continual learning strategies that can be applied to mitigate this issue, Memory replay (Rolnick et al., 2019; Aljundi et al., 2019a, b; Yoon et al., 2021) has been widely adopted due to its simplicity and effectiveness, especially when combined with a contrastive learning objective (Yoon et al., 2023c). These memory-based approaches typically maintain an external memory initialized with a set of queries and their corresponding relevant document pairs (i.e., positive samples). As new documents arrive, the encoder is updated to make each query closer to the positive samples while pushing it away from irrelevant documents (i.e., negative samples). Recent works such as L2R (Cai et al., 2023) and CLEVER (Chen et al., 2023) represent state-of-the-art frameworks for DR- and GR-based continual retrieval, respectively. L2R defines negativity and diversity metrics to mitigate false negatives during negative sampling. CLEVER leverages external memory for a pseudo-query generator with a reconstruction-based objective, a regularization-based objective, and dynamic indexing via Incremental Product Quantization (IPQ). However, both approaches present practical limitations in their supervision assumption: L2R requires a fixed set of queries along with ground-truth documents for training, and CLEVER also depends on a large number of positive query-document pairs for effective model initialization.
@@ 4.1. Retrieval
An effective retrieval system requires capturing the subtle contextual semantics between queries and documents, which is more pronounced in a label-free setting for continual retrieval. To minimize the loss of information from token embeddings of queries and documents, we aim to preserve token-level granularity in relevance estimation. Specifically, instead of relying on a single embedding derived from mean pooling or the [CLS] token, we adopt a token-level similarity approach inspired by ColBERT’s late interaction (Santhanam et al., 2022) to preserve the semantic granularity of individual tokens. Unlike ColBERT, however, we do not modify the encoder architecture or use special tokens to maintain simplicity and efficiency.
@@ 4.2. Memory Update
A benefit function is defined to trade off the gain in accuracy against the computational cost. The Johnson-Lindenstrauss (JL) lemma (Johnson et al., 1984) provides a cost model in terms of distortion $\varepsilon$, while the approximation quality imposes a lower bound on $\varepsilon$. Maximizing the benefit within this feasible range yields the optimal distortion rate $\varepsilon^{*}=\frac{1}{3\sqrt{e}}$. Using the JL lemma, we derive the minimal bit size for LSH at $\varepsilon^{*}$. See Appendix A.1 for the full proof. ∎
For example, in our evaluation setting with the LOTTE dataset (Santhanam et al., 2022), each session includes approximately 2,430 queries (with an average of 9 tokens) and 500,000 documents (with an average of 159 tokens), resulting in up to 80 million token embeddings per session, each with 768 dimensions by BERT (Devlin et al., 2019). Then, according to Theorem 4.1, to maintain an acceptable distortion rate of $\varepsilon=\frac{1}{3\sqrt{e}}\approx 0.2$, the sufficient number of RP-LSH bits is given by:
@@
CREAM summarizes each cluster by the compact triplet $\langle N,LS,SS\rangle$, similar to the cluster feature vector in BIRCH (Zhang et al., 1996). In CREAM, however, we track only the distance summaries, which are even more compact than the original embedding summaries in BIRCH. Specifically, $N$ denotes the number of instances in the cluster, $LS$ is the linear sum of distances to the prototype, and $SS$ is the sum of squared distances to the prototype. This cluster summary is sufficient to compute key cluster statistics, such as the mean and standard deviation of distances to the prototype, and also allows efficient incremental updates in an additive manner.
@@ 4.3. Encoder Update
To construct a training sample that effectively reflects the entire knowledge space in the memory, we propose a query sampling strategy by stratified sampling and coreset selection (Sener and Savarese, 2018). A relevant approach, topic-aware sampling (Hofstätter et al., 2021), has shown effectiveness in a static IR setting. While it randomly samples queries from a limited subset of clusters, CREAM considers the full cluster distribution and explicitly selects an optimal query set by maximizing coverage over the entire cluster space.
@@ 5.1. Experiment Setup
Datasets. To effectively model the dynamics of evolving data distributions, we conduct experiments on two real-world benchmark datasets widely used for continual retrieval (Cai et al., 2023; Gerald and Soulier, 2022): LoTTE (Santhanam et al., 2022) and MSMARCO (Lupart et al., 2023) with the queries clustered by topic from the original dataset (Nguyen et al., 2016). LoTTE includes five domains (writing, recreation, science, technology, and lifestyle) from StackExchange questions and answers. MSMARCO includes five domains (Names and Public Figures, Dated Events, Pricing/Units, Medical Treatments and Biology/Physics.) from Bing questions and answers. Each dataset is simulated in 10 streaming sessions with partially overlapping topics following the convention in continual learning (Wang et al., 2024; Buzzega et al., 2020). Based on the two evaluation protocols formalized in Section 3, the training and evaluation query sets are disjoint, and the training and evaluation document sets are either separated or shared. Details of datasets and sessions are given in Appendix A.4.
@@ A.5. Implementation Details
We use the BM25 [18] implementation from the Okapi library with $k_{1}=1.5$, $b=0.75$, and $\epsilon=0.25$. For ColBERT [20], we use the implementation provided in the official L2R codebase. Following the original paper, we set the output dimension of the linear projection layer in the model to 128. ColBERT+ performs incremental learning using negatives sampled from BM25-retrieved (but unannotated) documents. We use the official L2R implementations of Experience Replay (ER) [19], Maximally Interfered Retrieval (MIR) [1], and Gradient-based Sample Selection (GSS) [2]. Online Coreset Selection (OCS) [25] is implemented based on the L2R codebase, with $\alpha=1.0$, $\beta=1.0$, and $\gamma=1000.0$, following the original paper and code. For L2R [4], we use the official implementation with $\alpha=0.6$ and $\beta=0.4$, as configured in the code.
@@ A.8. LSH Bit Size for MSMARCO
Each session includes approximately 2,430 queries (with an average of 32 tokens) and 30,000 documents (with an average of 256 tokens), resulting in up to 8 million token embeddings per session, each with 768 dimensions by BERT [6]. Then, according to Theorem 4.1, to maintain an acceptable distortion rate of $\varepsilon=\frac{1}{3\sqrt{e}}\approx 0.2$, the minimum number of RP-LSH bits required is: